### v1

Abstract
In this paper, the light-front wave functions (LFWFs) of $\rho$ meson are evaluated in the framework of Nambu–Jona-Lasinio model using the proper time regularization scheme. The transverse momentum dependent parton distributions (TMDs) of the $\rho$ meson are derived from the overlap representations of the LFWFs. We investigate the $\bm{k}_{\perp}$-weighted moments and the $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ of $\rho$ meson TMDs, $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ shows the typical transverse momenta of quark TMDs in our model. Our findings indicate that the average transverse momenta of $\langle k_{\perp}(x)\rangle_{\alpha}$ fall within the range of $[0.3,0.45]$ GeV, while $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$ are in the region of $[0.15,0.30]$ GeV. Our TMDs and parton distribution function (PDFs) exhibit excellent adherence to positive constraints. The investigation of spin densities within the $\rho$ meson in transverse momentum space reveals axially symmetric distributions for quarks and targets polarized in either longitudinal or transverse directions. Conversely, unpolarized symmetric distributions occur when the quark is longitudinally polarized and the target is transversely polarized, while exhibiting dipolar distortions in the opposite scenario. A comparison between TMDs from LFWFs and those from a covariant approach demonstrates that TMDs from LFWFs more effectively satisfy positive constraints than their counterparts from the covariant approach. Furthermore, the $x$-moments of $\rho$ meson PDFs indicate that $\langle x\rangle_{g_{L}}$ is larger using LFWFs compared to the covariant approach. Additionally, analysis of quark spin densities within the $\rho$ meson reveals that $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)=f-g_{L}-\frac{1}{3}f_{LL}$ yields a negative result using the covariant approach, whereas $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)$ derived from LFWFs produces a positive outcome.
I Introduction
The most difficult job of hadron physics is to compute the hadronic transition matrix elements which are fully governed by non-perturbative Quantum Chromodynamics (QCD). Due to the lack of complete understanding of the non-perturbative QCD so far, one needs to utilize phenomenological models. In describing various processes, such models have achieved relative successes up to now. Among these models the Nambu–Jona-Lasinio (NJL) model Klevansky (1992); Buballa (2005); Cloët et al. (2014); Cui et al. (2017, 2016); Zhang et al. (2016, 2018); Zhang and Wu (2024) is a Poincaré covariant quantum field theory with many of the same low-energy properties as QCD. It has the same flavor symmetries as QCD and should therefore offer a stable chiral effective theory of QCD valid at low to intermediate energies. Above all, it includes the key prominent phenomena of confinement and dynamical chiral symmetry breaking.
The basic degrees of freedom of QCD are quarks and gluons, LFWFs Lepage and Brodsky (1980); Miller and Tiburzi (2010); Brodsky et al. (2004, 2010); Ke and Li (2011); Ji and Liu (2022); Zhang et al. (2021a); Zhang and Ping (2021); Lappi et al. (2020) serve as the interpolating functions that connect hadrons with their underlying quark and gluon constituents. LFWF is proved to be Lorentz invariant and expressed in terms of the momentum fractions of the constituents which are independent of the total hadron momentum. As a result, LFWFs play a crucial role in the light-front quantization approach to QCD and our understanding of hadron structure. Therefore, gaining insight into how to derive LFWFs from a fundamental perspective is highly beneficial.
A variety of high energy hadronic processes and experimentally observable quantities can be interpreted in terms of LFWFs, for example, electromagnetic form factors (FFs) Bartel et al. (1973); Hohler et al. (1976); Vogel and Engel (1989); Mergell et al. (1996); Kelly (2004); Perdrisat et al. (2007); Puckett et al. (2010); De Mori (2022); Stamen et al. (2022); Batelaan et al. (2022) and PDFs Holt and Roberts (2010); Nguyen et al. (2011); Dulat et al. (2016); Ding et al. (2020); Cui et al. (2021, 2020) can be analyzed using LFWFs. In addition, LFWFs also can be used to estimate of weak decay rates Brodsky and Hwang (1999); Choi and Ji (2007), calculate single spin asymmetries in semi-inclusive deep inelastic scattering Brodsky et al. (2002); Gamberg et al. (2008), and determine the cross sections of electromagnetic production for vector mesons Radyushkin (1996); Vanderhaeghen et al. (1999); Goloskokov and Kroll (2007). Furthermore, the three-dimensional structure of the nucleon in the momentum and coordinate space characterized by the TMDs Collins et al. (2008); Meissner et al. (2007); Hautmann (2007); Barone et al. (2010); Puhan and Dahiya (2024); Xie and Lu (2023); Puhan et al. (2024) and generalized parton distributions (GPDs) Müller et al. (1994); Ji (1997); Radyushkin (1997); Vanderhaeghen et al. (1999); Belitsky et al. (2002); Goeke et al. (2001); Burkardt (2003); Diehl (2003); Belitsky and Radyushkin (2005); Zhang et al. (2021b, c, 2022a); Adhikari et al. (2021); Fu et al. (2022); Zhang et al. (2022b); Luan and Lu (2024). TMDs are sensitive to spin dependent correlations and the parton’s intrinsic transverse motion, offering a more complete understanding of the parton structures, especially the transverse structures of hadrons.
Computing light vector meson LFWFs in QCD presents a significant challenge. The diagonalization of the light-cone QCD Hamiltonian becomes exceedingly challenging as the number of Fock-states Brodsky et al. (1998) increases, due to the numerous possibilities for the creation and annihilation of light-quarks and gluons. By calculating the hadronic eigensolutions $|\psi_{H}\rangle$ of the QCD light-front Hamiltonian $H_{LF}$ projected on the free Fick basis $\psi_{n}^{H}=\langle n|\psi_{H}\rangle$, one can obtain LFWFs. In principle, it is possible to solve for the hadronic LFWFs directly from fundamental theory by employing nonperturbative techniques such as discretized light-front quantization, the transverse lattice, lattice gauge theory moments, or Bethe-Salpeter (BSE)/Dyson-Schwinger equation (DSE) Shi et al. (2021, 2022). Reviews of nonperturbative light-front methods can be found in Refs. Carbonell et al. (1998); Brodsky et al. (1998); Dalley (2002).
In this paper, we have computed the leading Fock-state LFWFs of the $\rho$ meson within the framework of the Nambu-Jona-Lasinio (NJL) model. Specifically, we focus on the $\rho$ meson’s $|q\bar{q}\rangle$ component. The NJL model is characterized by an effective Lagrangian for relativistic fermions interacting through local fermion-fermion couplings, and it finds widespread application in various fields.
This paper is organized as follows: In Sec. II, we provide an introduction to the NJL model and then proceed to calculate the $\rho$ meson LFWFs within the framework of the NJL model. In Sec. III, we calculate the TMDs of the $\rho$ meson from the overlap representation of the LFWFs. The $\bm{k}_{\perp}$-weighted moments and the $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ of $\rho$ TMDs are studied. The positive bounds for $\rho$ meson TMDs and PDFs are checked. Additionally, we examine the spin densities of $\rho$ meson in transverse momentum space. Finally, a concise summary and outlook are provided in Section IV.
II $\rho$ meson LFWFs in the NJL model
II.1 NJL Model
The SU(2) flavor NJL Lagrangian is
$\mathcal{L}$ $=\bar{\psi}\left(i\gamma^{\mu}\partial_{\mu}-\hat{m}\right)\psi$
$+G_{\pi}\left[\left(\bar{\psi}\psi\right)^{2}-\left(\bar{\psi}\gamma_{5}\vec{\tau}\psi\right)^{2}\right]-G_{\omega}\left(\bar{\psi}\gamma_{\mu}\psi\right)^{2}$
$-G_{\rho}\left[\left(\bar{\psi}\gamma_{\mu}\vec{\tau}\psi\right)^{2}+\left(\bar{\psi}\gamma_{\mu}\gamma_{5}\vec{\tau}\psi\right)^{2}\right],$ (1)
where $\hat{m}=\text{diag}\left(m_{u},m_{d}\right)$ is current quark mass matrix and $\vec{\tau}$ are the Pauli matrices to stand for isospin. In the limit of exact isospin symmetry, $m_{u}=m_{d}=m$. The 4-fermion coupling constants in chiral channels are labeled by $G_{\pi}$ , $G_{\omega}$, and $G_{\rho}$.
The dressed quark propagator is,
[…]
The interaction kernel of the gap equation is local, one can obtain a constant dressed quark mass $M$, which satisfies
$M=m+12iG_{\pi}\int\frac{\mathrm{d}^{4}l}{(2\pi)^{4}}\text{tr}_{\text{D}}[S(l)],$ (3)
where the trace is over Dirac indices.
NJL model is not renormalizable, it needs a regularization scheme to fully define the model. We will use the proper time regularization scheme Ebert et al. (1996); Hellstern et al. (1997); Bentz and Thomas (2001).
$\frac{1}{X^{n}}$ $=\frac{1}{(n-1)!}\int_{0}^{\infty}\mathrm{d}\tau\,\tau^{n-1}e^{-\tau X}$
$\rightarrow\frac{1}{(n-1)!}\int_{1/\Lambda_{\text{UV}}^{2}}^{1/\Lambda_{\text{IR}}^{2}}\mathrm{d}\tau\,\tau^{n-1}e^{-\tau X}$ (4)
where $X$ stands for a product of propagators that have been combined using Feynman parametrization. We introduce the infrared cutoff $\Lambda_{\text{IR}}$ to mimic confinement because it is absent in the NJL model. The infrared cutoff should be of the order $\Lambda_{\text{QCD}}$ and we choose $\Lambda_{\text{IR}}=0.240$ GeV. The coupling strength $G_{\pi}$, the momentum cutoff $\Lambda_{\text{UV}}$ and the current quark mass $m$ are determined through the Gell-Mann–Oakes–Renner (GMOR) relation, $f_{\pi}^{2}m_{\pi}^{2}=-m\langle\bar{\psi}\psi\rangle$ and gap equation $M=m-2G_{\pi}\langle\bar{\psi}\psi\rangle$, where $\langle\bar{\psi}\psi\rangle$ is two-quark condensate derived from QCD sum rules. $m_{\pi}=0.140$ GeV is the physical pion mass, $f_{\pi}=0.092$ GeV is the pion decay constant, $m$ and $M$ are the current and the constituent quark mass, respectively.
The self-energy of the pseudoscalar bubble diagram is
$\Pi_{PP}(Q^{2})=-\frac{3}{2\pi^{2}}\int_{0}^{1}\mathrm{d}x\int_{1/\Lambda_{\text{UV}}^{2}}^{1/\Lambda_{\text{IR}}^{2}}\mathrm{d}\tau\,\frac{1}{\tau^{2}}e^{-\tau M^{2}}$
$+\frac{3}{4\pi^{2}}\int_{0}^{1}\mathrm{d}x\int_{1/\Lambda_{\text{UV}}^{2}}^{1/\Lambda_{\text{IR}}^{2}}\mathrm{d}\tau\,\frac{Q^{2}}{\tau}\,e^{-\tau(M^{2}+x\bar{x}Q^{2})},$ (5)
where $\bar{x}=1-x$. $m_{\pi}=0.14$ GeV is coinciding with the value obtained from the pole condition $01+2G_{\pi}\Pi_{PP}(-m_{\pi}^{2})=$ 0. $G_{\omega}$ and $G_{\rho}$ are determined by $m_{\omega}=0.782$ GeV, $m_{\rho}=0.770$ GeV through $01+2G_{i}\Pi_{VV}(-m_{i}^{2})=$ 0, $i=(\omega,\rho)$, where $\Pi_{VV}(Q^{2})$ is the self-energy of the vector bubble diagram defined in Eq. (7). The parameters used in this work are listed in Table 1.
From the inhomogeneous Bethe-Salpeter equation, the dressed quark form factors, associated with the electromagnetic currents, are Zhang et al. (2021a)
$F_{1i}(Q^{2})=\frac{1}{1+2G_{i}\Pi_{VV}(Q^{2})},\quad\quad F_{2i}(Q^{2})=0,$ (6)
where $i=(\omega,\rho)$, $\Pi_{VV}$ is the self-energy of the vector bubble diagram
$\Pi_{VV}(Q^{2})=\frac{3}{\pi^{2}}\int_{0}^{1}\mathrm{d}x\int_{1/\Lambda_{\text{UV}}^{2}}^{1/\Lambda_{\text{IR}}^{2}}\mathrm{d}\tau\,\frac{x\bar{x}}{\tau}Q^{2}e^{-\tau(M^{2}+x\bar{x}Q^{2})}.$ (7)
Table 1: Parameters used in our work. The dressed quark mass and regularization parameters are in units of GeV, while coupling constant is in units of GeV-2.
$\Lambda_{\text{IR}}$ $\Lambda_{\text{UV}}$ $M$ $m$ $m_{\rho}$ $G_{\pi}$ $G_{\omega}$ $G_{\rho}$ $N_{T}$ $N_{L}$
0.24 0.645 0.4 0.016 0.77 19.0 10.4 11.0 3.618 3.005
II.2 The definition and calculation of LFWFs
Considering dynamical spin effects, the vector meson wave functions in the NJL model can be written as
$\Psi_{\lambda,\lambda^{\prime}}^{\Lambda}(x,\bm{k}_{\perp})=iN_{\Lambda}\int\frac{\mathrm{d}k^{+}\mathrm{d}k^{-}}{2\pi}\delta(x-\frac{k^{+}}{q^{+}})$
×$\times\frac{\sqrt{x\bar{x}}}{\left(k^{2}-M^{2}\right)\left((k-q)^{2}-M^{2}\right)}\chi_{\lambda,\lambda^{\prime}}^{\Lambda}(x,\bm{k}_{\perp}),$ (8)
where $a^{0}=\frac{1}{\sqrt{2}}(a^{0}+a^{3})$, $\bm{a}_{\perp}=(a_{1},a_{2})$ are the light-cone $\pm$ and transverse components of a 4-vector $a^{\mu}$. The Lorentz invariant spin structure for the vector meson is expressed by accounting the photon-quark-antiquark vertex:
$\chi_{\lambda,\lambda^{\prime}}^{\Lambda}(x,\bm{k}_{\perp})$ $=\frac{\bar{u}_{\lambda}(xq^{+},\bm{k}_{\perp})}{\sqrt{x}}\epsilon_{\Lambda}\cdot\gamma\frac{v_{\lambda^{\prime}}(\bar{x}q^{+},-\bm{k}_{\perp})}{\sqrt{1-x}},$ (9)
where $xq^{+}$ and $\bar{x}q^{+}$ are the longitudinal momentum of the quark and antiquark, $\bm{k}_{\perp}$ and $-\bm{k}_{\perp}$ are the transverse momentum of the quark and qntiquark respectively. $\epsilon_{\Lambda}$ are the polarizations vectors, for the longitudinally polarized and the transversely polarized $\rho$ meson are given by
$\epsilon_{L}=(\frac{q^{+}}{m_{\rho}},-\frac{m_{\rho}}{q^{+}},0,0),\quad\quad\epsilon_{T}^{\pm}=\mp\frac{1}{\sqrt{2}}(0,0,1,\pm i),$ (10)
then we obtain the final form of
$\Psi_{\lambda,\lambda^{\prime}}^{0}$ $=\frac{\sqrt{6}N_{L}(x\bar{x}m_{\rho}^{2}+M^{2}+\bm{k}_{\perp}^{2})\delta_{\lambda,-\lambda^{\prime}}}{(\bm{k}_{\bot}^{2}+M^{2}-\bar{x}xm_{\rho}^{2})},$ (11)
$\Psi_{\lambda,\lambda^{\prime}}^{+1}$ $=\frac{\sqrt{6}N_{T}\left(\delta_{\lambda+,\lambda^{\prime}-}xk_{\lambda}-k_{-\lambda}\bar{x}\delta_{\lambda-,\lambda^{\prime}+}+M\delta_{\lambda+,\lambda^{\prime}+}\right)}{\bm{k}_{\bot}^{2}+M^{2}-\bar{x}xm_{\rho}^{2}},$ (12)
$\Psi_{\lambda,\lambda^{\prime}}^{-1}$ $=\frac{\sqrt{6}N_{T}\left(\delta_{\lambda-,\lambda^{\prime}+}xk_{\lambda}-k_{-\lambda}\bar{x}\delta_{\lambda+,\lambda^{\prime}-}-M\delta_{\lambda-,\lambda^{\prime}-}\right)}{\bm{k}_{\bot}^{2}+M^{2}-\bar{x}xm_{\rho}^{2}},$ (13)
where $k_{\lambda}=k_{i}+i\lambda k_{2}$, $\delta_{ab,cd}=\delta_{a,b}\delta_{c,d}$, with $\delta_{a,b}$ the Kronecker delta. The LFWFs are divergent, we should apply the PTR to make them convergent. $\Psi_{\pm,\pm}^{\Lambda=0}$ and $\Psi_{\pm,\pm}^{\Lambda=\pm 1}$ represent $|l_{z}|=1$ and $|l_{z}|=2$ respectively.
The normalization constants $N_{T(L)}$ are determined by
$\sum_{\lambda,\lambda^{\prime}}\int\frac{\mathrm{d}x\mathrm{d}^{2}\bm{k}_{\perp}}{2(2\pi)^{3}}|\Psi_{\lambda,\lambda^{\prime}}^{\Lambda}|^{2}=1,$ (14)
depending upon the polarization of the $\rho$ meson. We obtain $N_{T}^{2}=13.092$ and $N_{L}^{2}=9.032$.
III $\rho$ meson TMDs
The transverse momentum-dependent quark correlation function of spin-1 particle is defined as:
$\bm{\Phi}_{\beta\alpha}^{(\lambda)\bm{S}}(x,\bm{k}_{\perp})$
$=\int\frac{\mathrm{d}k^{+}\mathrm{d}k^{-}}{(2\pi)^{4}}\delta(k^{+}-xp^{+})$
×$\times\int\mathrm{d}^{4}ze^{ik\cdot z}\,_{\bm{S}}\langle p,\lambda|\bar{\psi}_{\alpha}(0)\psi_{\beta}(z)|p,\lambda\rangle_{\bm{S}},$
$=\int\frac{\mathrm{d}z^{-}\mathrm{d}^{2}\bm{z}_{\perp}}{(2\pi)^{3}}\,_{\bm{S}}\langle p,\lambda|\bar{\psi}_{\alpha}(0)\psi_{\beta}(z^{-},\bm{z}_{\perp})|p,\lambda\rangle_{\bm{S}}$
×$\times e^{i(xp^{+}z^{-}-\bm{k}_{\perp}\cdot\bm{z}_{\perp})}$
$=\varepsilon_{(\lambda)\mu}^{*}(p)\bm{\Phi}_{\beta\alpha}^{\mu\nu}(x,\bm{k}_{\perp})\varepsilon_{(\lambda)\nu}(p),$ (15)
where $\psi$ stands for the flavor SU(2) quark field operator, $\alpha$ and $\beta$ are Dirac indices, for any four-vector. The 3-momentum of the target ($\bm{p}$) is assumed to be in the $z$-direction, and the quark’s momentum components perpendicular to this direction are denoted by $\bm{k}_{\perp}$. The state $|p,\lambda\rangle$ denotes that the projection of the target’s spin on the direction $\bm{S}$ is equal to $\lambda=\pm 1$, 00, and is normalized according to the covariant normalization $\langle p,\lambda^{{}^{\prime}}|p,\lambda\rangle=2p^{+}V\delta_{\lambda,\lambda^{{}^{\prime}}}$, where $V$ represents the quantization volume.
At leading-twist, the coefficient functions are represented by
$\langle\gamma^{+}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp})=\frac{1}{2}\text{Tr}_{D}\left[\gamma^{+}\Phi^{(\lambda)\bm{S}}(x,\bm{k}_{\perp})\right]$
$\equiv\varepsilon_{(\lambda)\mu}^{*}(p)\langle\gamma^{+}\rangle^{\mu\nu}(x,\bm{k}_{\perp})\varepsilon_{(\lambda)\nu}(p),$ (16)
$\langle\gamma^{+}\gamma_{5}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp})=\frac{1}{2}\text{Tr}_{D}\left[\gamma^{+}\gamma^{5}\Phi^{(\lambda)\bm{S}}(x,\bm{k}_{\perp})\right]$
$\equiv\varepsilon_{(\lambda)\mu}^{*}(p)\langle\gamma^{+}\gamma_{5}\rangle^{\mu\nu}(x,\bm{k}_{\perp})\varepsilon_{(\lambda)\nu}(p),$ (17)
$\langle\gamma^{+}\gamma^{i}\gamma^{5}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp})=\frac{1}{2}\text{Tr}_{D}\left[\gamma^{+}\gamma^{i}\gamma^{5}\Phi^{(\lambda)\bm{S}}(x,\bm{k}_{\perp})\right]$
$\equiv\varepsilon_{(\lambda)\mu}^{*}(p)\langle\gamma^{+}\gamma^{i}\gamma^{5}\rangle^{\mu\nu}(x,\bm{k}_{\perp})\varepsilon_{(\lambda)\nu}(p),$ (18)
where $i=1,2$ represents the transverse vector components. According to Refs. Bacchetta and Mulders (2001); Ninomiya et al. (2017), the coefficient functions of the spin-1 hadron TMDs can be expressed as follows:
$\langle\gamma^{+}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})\equiv f_{1}(x,\bm{k}_{\perp}^{2})+S_{LL}f_{LL}(x,\bm{k}_{\perp}^{2})$
$+\frac{\bm{S}_{LT}\cdot\bm{k}_{\perp}}{m_{\rho}}f_{LT}(x,\bm{k}_{\perp}^{2})+\frac{\bm{k}_{\perp}\cdot\bm{S}_{TT}\cdot\bm{k}_{\perp}}{m_{\rho}^{2}}f_{TT}(x,\bm{k}_{\perp}^{2}),$ (19)
$\langle\gamma^{+}\gamma^{5}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})$ $\equiv\lambda\left[S_{L}g_{L}(x,\bm{k}_{\perp}^{2})+\frac{\bm{S}_{T}\cdot\bm{k}_{\perp}}{m_{\rho}}g_{T}(x,\bm{k}_{\perp}^{2})\right],$ (20)
$\langle\gamma^{+}\gamma^{i}\gamma^{5}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})\equiv\lambda\left[S_{T}^{i}h(x,\bm{k}_{\perp}^{2})+S_{L}\frac{k_{\perp}^{i}}{m_{\rho}}h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})\right]$
$+\frac{\lambda}{2m_{\rho}^{2}}(2k_{\perp}^{i}\bm{k}_{\perp}\cdot\bm{S}_{T}-S_{T}^{i}\bm{k}_{\perp}^{2})h_{T}^{\perp}(x,\bm{k}_{\perp}^{2}),$ (21)
where the specific $\bm{S}$ and $\lambda$ dependence are:
$S_{LL}$ $=(3\lambda^{2}-2)(\frac{1}{6}-\frac{1}{2}S_{L}^{2})\,,$ (22a)
$S_{LT}^{i}$ $=(3\lambda^{2}-2)S_{L}S_{T}^{i}\,,$ (22b)
$S_{TT}^{ij}$ $=(3\lambda^{2}-2)(S_{T}^{i}S_{T}^{j}-\frac{1}{2}\bm{S}_{T}^{2}\delta^{ij}).$ (22c)
$f_{LL}(x,\bm{k}_{\perp}^{2})$, $f_{LT}(x,\bm{k}_{\perp}^{2})$ and $f_{TT}(x,\bm{k}_{\perp}^{2})$ are specific to tensor polarized hadrons with spin $J\geq 1$.
III.1 Overlap formalism
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 1: Results for the quark TMDs of $\rho$ meson as a function of $x$ and $\bm{k}_{\perp}^{2}$. The TMDs are given in units of GeV-2 and $\bm{k}_{\perp}^{2}$ in units of GeV2.
The explicit overlap form for all the T-even TMDs in terms of helicity amplitudes is expressed as Shi et al. (2022)
$f(x,\bm{k}_{\perp}^{2})$ $=\frac{1}{6(2\pi)^{3}}\sum_{h_{q},h_{\bar{q}}}(|\Psi_{h_{q},h_{\bar{q}}}^{0}|^{2}+|\Psi_{h_{q},h_{\bar{q}}}^{+1}|^{2}+|\Psi_{h_{q},h_{\bar{q}}}^{-1}|^{2})\,,$ (23a)
$g_{L}(x,\bm{k}_{\perp}^{2})$ $=\frac{1}{4(2\pi)^{3}}\sum_{h_{\bar{q}}}(|\Psi_{+,h_{\bar{q}}}^{+1}|^{2}-|\Psi_{-,h_{\bar{q}}}^{+1}|^{2}-|\Psi_{+,h_{\bar{q}}}^{-1}|^{2}+|\Psi_{-,h_{\bar{q}}}^{-1}|^{2})\,,$ (23b)
$g_{T}(x,\bm{k}_{\perp}^{2})$ $=\frac{m_{\rho}}{4\sqrt{2}(2\pi)^{3}\bm{k}_{\perp}^{2}}\sum_{h_{\bar{q}}}(k_{R}(\Psi_{+,h_{\bar{q}}}^{+1*}\Psi_{+,h_{\bar{q}}}^{0}-\Psi_{-,h_{\bar{q}}}^{+1*}\Psi_{-,h_{\bar{q}}}^{0}+\Psi_{+,h_{\bar{q}}}^{0*}\Psi_{+,h_{\bar{q}}}^{-1}-\Psi_{-,h_{\bar{q}}}^{0*}\Psi_{-,h_{\bar{q}}}^{-1}))$
$+\frac{m_{\rho}}{4\sqrt{2}(2\pi)^{3}\bm{k}_{\perp}^{2}}\sum_{h_{\bar{q}}}(k_{L}(\Psi_{+,h_{\bar{q}}}^{0*}\Psi_{+,h_{\bar{q}}}^{+1}-\Psi_{-,h_{\bar{q}}}^{0*}\Psi_{-,h_{\bar{q}}}^{+1}+\Psi_{+,h_{\bar{q}}}^{-1*}\Psi_{+,h_{\bar{q}}}^{0}-\Psi_{-,h_{\bar{q}}}^{-1*}\Psi_{-,h_{\bar{q}}}^{0}))\,,$ (23c)
$h(x,\bm{k}_{\perp}^{2})$ $=\frac{1}{4\sqrt{2}(2\pi)^{3}}\sum_{h_{\bar{q}}}(\Psi_{+,h_{\bar{q}}}^{+1*}\Psi_{-,h_{\bar{q}}}^{0}+\Psi_{-,h_{\bar{q}}}^{0*}\Psi_{+,h_{\bar{q}}}^{+1}+\Psi_{+,h_{\bar{q}}}^{0*}\Psi_{-,h_{\bar{q}}}^{-1}+\Psi_{-,h_{\bar{q}}}^{-1*}\Psi_{+,h_{\bar{q}}}^{0})\,,$ (23d)
$h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})$ $=\frac{m_{\rho}}{4(2\pi)^{3}\bm{k}_{\perp}^{2}}\sum_{h_{\bar{q}}}(k_{R}(\Psi_{-,h_{\bar{q}}}^{+1*}\Psi_{+,h_{\bar{q}}}^{+1}-\Psi_{-,h_{\bar{q}}}^{-1*}\Psi_{+,h_{\bar{q}}}^{-1})+k_{L}(\Psi_{+,h_{\bar{q}}}^{+1*}\Psi_{-,h_{\bar{q}}}^{+1}-\Psi_{+,h_{\bar{q}}}^{-1*}\Psi_{-,h_{\bar{q}}}^{-1}))\,,$ (23e)
$h_{T}^{\perp}(x,\bm{k}_{\perp}^{2})$ $=\frac{m_{\rho}^{2}}{2\sqrt{2}(2\pi)^{3}\bm{k}_{\perp}^{4}}\sum_{h_{\bar{q}}}(k_{R}^{2}(\Psi_{-,h_{\bar{q}}}^{+1*}\Psi_{+,h_{\bar{q}}}^{0}+\Psi_{-,h_{\bar{q}}}^{0*}\Psi_{+,h_{\bar{q}}}^{-1})+k_{L}^{2}(\Psi_{+,h_{\bar{q}}}^{0*}\Psi_{-,h_{\bar{q}}}^{+1}+\Psi_{+,h_{\bar{q}}}^{-1*}\Psi_{-,h_{\bar{q}}}^{0}))\,,$ (23f)
$f_{LL}(x,\bm{k}_{\perp}^{2})$ $=\frac{1}{2(2\pi)^{3}}\sum_{h_{q},h_{\bar{q}}}(|\Psi_{h_{q},h_{\bar{q}}}^{0}|^{2}-\frac{1}{2}(|\Psi_{h_{q},h_{\bar{q}}}^{+1}|^{2}+|\Psi_{h_{q},h_{\bar{q}}}^{-1}|^{2}))\,,$ (23g)
$f_{LT}(x,\bm{k}_{\perp}^{2})$ $=\frac{m_{\rho}}{4\sqrt{2}(2\pi)^{3}\bm{k}_{\perp}^{2}}\sum_{h_{\bar{q}}}(k_{R}(\Psi_{+,h_{\bar{q}}}^{+1*}\Psi_{+,h_{\bar{q}}}^{0}+\Psi_{-,h_{\bar{q}}}^{+1*}\Psi_{-,h_{\bar{q}}}^{0}-\Psi_{+,h_{\bar{q}}}^{0*}\Psi_{+,h_{\bar{q}}}^{-1}-\Psi_{-,h_{\bar{q}}}^{0*}\Psi_{-,h_{\bar{q}}}^{-1}))$
$+\frac{m_{\rho}}{4\sqrt{2}(2\pi)^{3}\bm{k}_{\perp}^{2}}\sum_{h_{\bar{q}}}(k_{L}(\Psi_{+,h_{\bar{q}}}^{0*}\Psi_{+,h_{\bar{q}}}^{+1}+\Psi_{-,h_{\bar{q}}}^{0*}\Psi_{-,h_{\bar{q}}}^{+1}-\Psi_{+,h_{\bar{q}}}^{-1*}\Psi_{+,h_{\bar{q}}}^{0}-\Psi_{-,h_{\bar{q}}}^{-1*}\Psi_{-,h_{\bar{q}}}^{0}))\,,$ (23h)
$f_{TT}(x,\bm{k}_{\perp}^{2})$ $=\frac{m_{\rho}^{2}}{4(2\pi)^{3}\bm{k}_{\perp}^{4}}\sum_{h_{\bar{q}}}(k_{R}^{2}(\Psi_{+,h_{\bar{q}}}^{+1*}\Psi_{+,h_{\bar{q}}}^{-1}+\Psi_{-,h_{\bar{q}}}^{+1*}\Psi_{-,h_{\bar{q}}}^{-1})+k_{L}^{2}(\Psi_{+,h_{\bar{q}}}^{-1*}\Psi_{+,h_{\bar{q}}}^{+1}+\Psi_{-,h_{\bar{q}}}^{-1*}\Psi_{-,h_{\bar{q}}}^{+1}))\,,$ (23i)
where $k_{R(L)}=k_{1}\pm ik_{2}$, using the LFWFs in Eqs. (11) - (13), we can derive explicit expressions for the leading-twist T-even TMDs for the $\rho$ meson
$f(x,\bm{k}_{\perp}^{2})$ $=\frac{2N_{L}^{2}}{(2\pi)^{3}}\frac{\left(x\bar{x}m_{\rho}^{2}+M^{2}+\bm{k}_{\perp}^{2}\right)^{2}}{[A]^{2}}$
$+\frac{2N_{T}^{2}}{(2\pi)^{3}}\frac{(M^{2}+x^{2}\bm{k}_{\bot}^{2}+\bar{x}^{2}\bm{k}_{\bot}^{2})}{[A]^{2}}\,,$ (24a)
$g_{L}(x,\bm{k}_{\perp}^{2})$ $=\frac{3N_{T}^{2}}{(2\pi)^{3}}\frac{(M^{2}+(2x-1)\bm{k}_{\bot}^{2})}{[A]^{2}}\,,$ (24b)
$g_{T}(x,\bm{k}_{\perp}^{2})$ $=\frac{6N_{T}N_{L}m_{\rho}}{\sqrt{2}(2\pi)^{3}}\frac{\left(x\bar{x}m_{\rho}^{2}+M^{2}+\bm{k}_{\perp}^{2}\right)}{[A]^{2}}\,,$ (24c)
$h(x,\bm{k}_{\perp}^{2})$ $=\frac{6N_{T}N_{L}M}{\sqrt{2}(2\pi)^{3}}\frac{\left(x\bar{x}m_{\rho}^{2}+M^{2}+\bm{k}_{\perp}^{2}\right)}{[A]^{2}}\,,$ (24d)
$h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})$ $=-\frac{6N_{T}^{2}}{(2\pi)^{3}}\frac{\bar{x}Mm_{\rho}}{[A]^{2}}\,,$ (24e)
$h_{T}^{\perp}(x,\bm{k}_{\perp}^{2})$ $=0\,,$ (24f)
$f_{LL}(x,\bm{k}_{\perp}^{2})$ $=\frac{6N_{L}^{2}}{(2\pi)^{3}}\frac{\left(x\bar{x}m_{\rho}^{2}+M^{2}+\bm{k}_{\perp}^{2}\right)^{2}}{[A]^{2}}$
$-\frac{3N_{T}^{2}}{(2\pi)^{3}}\frac{(M^{2}+x^{2}\bm{k}_{\bot}^{2}+\bar{x}^{2}\bm{k}_{\bot}^{2})}{[A]^{2}}\,,$ (24g)
$f_{LT}(x,\bm{k}_{\perp}^{2})$ $=\frac{6N_{T}N_{L}m_{\rho}}{\sqrt{2}(2\pi)^{3}}\frac{(2x-1)\left(x\bar{x}m_{\rho}^{2}+M^{2}+\bm{k}_{\perp}^{2}\right)}{[A]^{2}}\,,$ (24h)
$f_{TT}(x,\bm{k}_{\perp}^{2})$ $=\frac{6N_{T}^{2}m_{\rho}^{2}}{(2\pi)^{3}}\frac{x\bar{x}}{[A]^{2}},$ (24i)
where $A=\bm{k}_{\perp}^{2}+M^{2}+x(x-1)m_{\rho}^{2}$, these TMDs are divergent, apply the PTR scheme one can obtain the final form of $\rho$ meson TMDs,
$f$ $=\frac{N_{L}^{2}}{4\pi^{3}}\int\mathrm{d}\tau(3x\bar{x}m_{\rho}^{2}+M^{2}+\frac{1}{\tau}+4\tau x^{2}\bar{x}^{2}m_{\rho}^{4})e^{-\tau A}$
$+\frac{N_{T}^{2}}{4\pi^{3}}\int\mathrm{d}\tau(x^{2}+\bar{x}^{2})(1+\tau x\bar{x}m_{\rho}^{2})e^{-\tau A}$
$+\frac{N_{T}^{2}}{4\pi^{3}}\int\mathrm{d}\tau\tau(1+2x\bar{x}M^{2})e^{-\tau A}\,,$ (25a)
$g_{L}$ $=\frac{3N_{T}^{2}}{8\pi^{3}}\int\mathrm{d}\tau e^{-\tau A}$
×$\times((2x-1)(1-\tau(M^{2}-m_{\rho}^{2}\bar{x}x))+\tau M^{2})\,,$ (25b)
$g_{T}$ $=\frac{3N_{T}N_{L}}{4\sqrt{2}\pi^{3}}\int\mathrm{d}\tau m_{\rho}(1+2\tau x\bar{x}m_{\rho}^{2})e^{-\tau A}\,,$ (25c)
$h$ $=\frac{3N_{T}N_{L}M}{4\sqrt{2}\pi^{3}}\int\mathrm{d}\tau(1+2\tau x\bar{x}m_{\rho}^{2})e^{-\tau A}\,,$ (25d)
$h_{L}^{\perp}$ $=-\frac{3N_{T}^{2}}{4\pi^{3}}\int\mathrm{d}\tau\tau\bar{x}Mm_{\rho}e^{-\tau A}\,,$ (25e)
$h_{T}^{\perp}$ $=0\,,$ (25f)
$f_{LL}$ $=\frac{3N_{L}^{2}}{4\pi^{3}}\int\mathrm{d}\tau(3x\bar{x}m_{\rho}^{2}+M^{2}+\frac{1}{\tau}+4\tau x^{2}\bar{x}^{2}m_{\rho}^{4})e^{-\tau A}$
$-\frac{3N_{T}^{2}}{8\pi^{3}}\int\mathrm{d}\tau(x^{2}+\bar{x}^{2})(1+\tau x\bar{x}m_{\rho}^{2})e^{-\tau A}$
$-\frac{3N_{T}^{2}}{8\pi^{3}}\int\mathrm{d}\tau\tau(1+2x\bar{x}M^{2})e^{-\tau A}\,,$ (25g)
$f_{LT}$ $=\frac{3N_{T}N_{L}m_{\rho}}{4\sqrt{2}\pi^{3}}\int\mathrm{d}\tau(2x-1)(1+2\tau x\bar{x}m_{\rho}^{2})e^{-\tau A}\,,$ (25h)
$f_{TT}$ $=\frac{3N_{T}^{2}}{4\pi^{3}}\int\mathrm{d}\tau\tau x\bar{x}m_{\rho}^{2}e^{-\tau A}.$ (25i)
The three-dimensional structure of the eight non-zero valance quark TMDs in the NJL model is depicted in Fig. 1. From the diagram, , it can be observed that the TMDs exhibit symmetry around $x=1/2$, with the exception of $h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})$ and $g_{L}(x,\bm{k}_{\perp}^{2})$. For $h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})$ the minimum value occurs at approximately $x\simeq 0.4$ at $\bm{k}_{\perp}^{2}\simeq 0$, which is smaller than $0.5$, consistent with Refs. Kaur et al. (2021); Ninomiya et al. (2017). As $\bm{k}_{\perp}^{2}$ increases, the minimum value shifts to smaller values of $x$. Regarding $g_{L}(x,\bm{k}_{\perp}^{2})$, the maximum value is around $x\simeq 0.6$ for $\bm{k}_{\perp}^{2}\simeq 0$, where $x>1/2$. In Ref. Ninomiya et al. (2017), they also calculated the $\rho$ meson TMDs in the NJL model using a covariant approach by directly calculating them from Feynman diagrams in the NJL model. Our method involves first calculating the LFWFs of $\rho$ meson and then obtaining its TMDs from their overlap. The results of TMDs are similar except for $g_{L}(x,\bm{k}_{\perp}^{2})$. Our calculation yields positive values for all regions of $g_{L}(x,\bm{k}_{\perp}^{2})$, whereas in Ref. Ninomiya et al. (2017), $g_{L}(x,\bm{k}_{\perp}^{2})$ has negative values. This can also be observed in the PDFs, and we will discuss this later.
Similar to Ref. Ninomiya et al. (2017), our TMDs results in Eq. (25) show that
$g_{T}(x,\bm{k}_{\perp}^{2})=\frac{m_{\rho}}{M}h(x,\bm{k}_{\perp}^{2}),$ (26)
this relationship would not maintain QCD evolution. $h(x,\bm{k}_{\perp}^{2})$ and $h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})$ are proportional to dressed quark mass $M$.
III.2 Transverse momentum dependence
The most intriguing aspect of the TMDs is, undoubtedly, their dependence on transverse momentum. Our initial investigation focuses on the typical transverse momenta of unpolarized quarks in the NJL model. This dependence on transverse momentum of the TMDs can be further elucidated through $\bm{k}_{\perp}$-weighted moments.
$\langle k_{\perp}^{n}\rangle_{\alpha}$ $\equiv\frac{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\bm{k}_{\perp}|\bm{k}_{\perp}|^{n}\alpha(x,\bm{k}_{\perp}^{2})}{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\bm{k}_{\perp}\alpha(x,\bm{k}_{\perp}^{2})},$ (27)
where $n$ represents the order of the moment, $\alpha(x,\bm{k}_{\perp}^{2})$ is an arbitrary TMD. The results of $\langle k_{\perp}^{n}\rangle$ are listed in Table 2, where we also include the NJL results from Ref. Ninomiya et al. (2017). It can be observed from Table 2 that the numerical results are similar, except for $g_{L}$, as our results are larger than theirs. The $\bm{k}_{\perp}$-weighted moments of $g_{L}$ coincide with those of $h_{L}^{\perp}$, which is consistent with the findings in Refs. Kaur et al. (2021); Shi et al. (2022). In our results, both $f_{LL}(x)$ and $f_{LT}(x)$ are zero, leading to a vanishing denominator in the above equation.
Table 2: The $\langle k_{\perp}\rangle$, $\langle k_{\perp}^{2}\rangle$ and $\langle k_{\perp}^{3}\rangle$ moments of the various TMDs defined by Eq. (27), in units of GeV, GeV2 and GeV3 respectively. Note, the denominator of Eq. (27) vanishes for $f_{LL}(x,\bm{k}_{\perp}^{2})$ and $f_{LT}(x,\bm{k}_{\perp}^{2})$, so they are not listed.
$\langle k_{\perp}^{n}\rangle$ $f$ $g_{L}$ $g_{T}$ $h$ $h_{L}^{\perp}$ $f_{TT}$
Ref. Ninomiya et al. (2017) $\langle k_{\perp}\rangle$ 0.32 0.08 0.34 0.34 0.33 0.32
Ref. Ninomiya et al. (2017) $\langle k_{\perp}^{2}\rangle$ 0.13 –0.11 0.16 0.16 0.15 0.14
Our $\langle k_{\perp}\rangle$ 0.368 0.327 0.346 0.346 0.327 0.322
Our $\langle k_{\perp}^{2}\rangle$ 0.189 0.147 0.165 0.165 0.147 0.143
Our $\langle k_{\perp}^{3}\rangle$ 0.124 0.084 0.100 0.100 0.084 0.080
In order to obtain the typical transverse momenta of unpolarized quarks in the NJL model, we also introduce an $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle$, as described in Ref. Avakian et al. (2010)
$\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ $=\frac{\int\mathrm{d}^{2}\bm{k}_{\perp}|\bm{k}_{\perp}|^{n}\alpha(x,\bm{k}_{\perp}^{2})}{\int\mathrm{d}^{2}\bm{k}_{\perp}\alpha(x,\bm{k}_{\perp}^{2})},$ (28)
where $\alpha$ is quark TMDs. The $x$-dependent average transverse momentum $\langle k_{\perp}(x)\rangle_{f}$ and square $\langle k_{\perp}^{2}(x)\rangle_{f}$ are plotted in Fig. 2.
Figure 2: The $x$-dependent average transverse momentum and square: $\langle k_{\perp}(x)\rangle_{f}$ – solid red curve; $\langle k_{\perp}^{2}(x)\rangle_{f}$ – dashed blue curve.
We find that in the valence-$x$ region $\langle k_{\perp}(x)\rangle_{f}$ and $\langle k_{\perp}^{2}(x)\rangle_{f}$ depend on $x$ weakly. It has a minimum value around $x=1/2$ and symmetric around $x=1/2$, the maximum values are around $x=0$ and $x=1$. Numerically we find
$\langle k_{\perp}(x)\rangle_{f}$ $\approx 0.4\quad\text{GeV}\,,$ (29)
$\langle k_{\perp}^{2}(x)\rangle_{f}$ $\approx 0.25\quad\text{GeV},$ (30)
this weak dependence is similar to the findings in Ref. Avakian et al. (2010) regarding the $u$ quark TMDs of the nucleon, but the numerical values are larger than their results at low hadronic scales. At higher scales, larger values are required as indicated by previous studies D’Alesio and Murgia (2004); Anselmino et al. (2005); Collins et al. (2006), such as $0.64$ GeV from the EMC data in Ref. Anselmino et al. (2005) and $0.56$ GeV from HERMES data in Ref. Collins et al. (2006) for the $u$ quark’s $x$-dependent average transverse momentum $\langle k_{\perp}(x)\rangle_{f}$ of the nucleon.
Figure 3: Upper panel: the $x$-dependent average transverse momentum $\langle k_{\perp}(x)\rangle_{\alpha}$. $\langle k_{\perp}(x)\rangle_{f_{TT}}$ – the dashed blue curve; $\langle k_{\perp}(x)\rangle_{g_{T}}$ – the thin solid black curve; $\langle k_{\perp}(x)\rangle_{g_{L}}$ – the thick solid orange curve; $\langle k_{\perp}(x)\rangle_{h}$ – the dot-dashed purple curve; $\langle k_{\perp}(x)\rangle_{h_{L}^{\perp}}$ – the dotted green curve; and Lower panel: the $x$-dependent average transverse momentum square $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$. $\langle k_{\perp}^{2}(x)\rangle_{f_{TT}}$ – the dashed blue curve; $\langle k_{\perp}^{2}(x)\rangle_{g_{T}}$ – the thin solid black curve; $\langle k_{\perp}^{2}(x)\rangle_{g_{L}}$ – the thick solid orange curve; $\langle k_{\perp}^{2}(x)\rangle_{h}$ – the dot-dashed purple curve; $\langle k_{\perp}^{2}(x)\rangle_{h_{L}^{\perp}}$ – the dotted green curve.
The $x$-dependent average transverse momentum $\langle k_{\perp}(x)\rangle_{\alpha}$ and square $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$ of the remanent TMDs are depicted in Fig. 3. The plots illustrate that $\langle k_{\perp}^{1,2}(x)\rangle_{g_{T}}$ and $\langle k_{\perp}^{1,2}(x)\rangle_{h}$ coincide, and they are numerically the largest. Numerically, it is found that
$\langle k_{\perp}(x)\rangle_{g_{T}}$ $=\langle k_{\perp}(x)\rangle_{h}\simeq[0.35,0.42]\quad\text{GeV}\,,$ (31)
$\langle k_{\perp}^{2}(x)\rangle_{g_{T}}$ $=\langle k_{\perp}^{2}(x)\rangle_{h}\simeq[0.16,0.24]\quad\text{GeV},$ (32)
$\langle k_{\perp}^{1,2}(x)\rangle_{f_{TT}}$ and $\langle k_{\perp}^{1,2}(x)\rangle_{h_{L}^{\perp}}$ are also coincide with each other. Their behaviors are also similar, exhibiting a minimum value around $x=1/2$, symmetry around $x=1/2$, and maximum values at approximately $x=0$ and $x=1$. This similarity is further supported by Table 2, where it is shown that $\langle k_{\perp}^{n}\rangle_{g_{T}}=\langle k_{\perp}^{n}\rangle_{h}$. Numerically we find
$\langle k_{\perp}(x)\rangle_{f_{TT}}$ $=\langle k_{\perp}(x)\rangle_{h_{L}^{\perp}}\simeq[0.32,0.38]\quad\text{GeV}\,,$ (33)
$\langle k_{\perp}^{2}(x)\rangle_{f_{TT}}$ $=\langle k_{\perp}^{2}(x)\rangle_{h_{L}^{\perp}}\simeq[0.14,0.19]\quad\text{GeV}.$ (34)
Table 2 demonstrates that $\langle k_{\perp}^{n}\rangle_{f_{TT}}$ is not equal to $\langle k_{\perp}^{n}\rangle_{h_{L}^{\perp}}$. However, based on Fig. 3, we should have
$\int_{0}^{1}\mathrm{d}x\langle k_{\perp}^{n}(x)\rangle_{f_{TT}}=\int_{0}^{1}\mathrm{d}x\langle k_{\perp}^{n}(x)\rangle_{h_{L}^{\perp}}=0.332,$ (35)
which is slightly larger than $\langle k_{\perp}^{n}\rangle_{f_{TT}}$ and $\langle k_{\perp}^{n}\rangle_{h_{L}^{\perp}}$, this indicates that
$\int_{0}^{1}\mathrm{d}x\langle k_{\perp}^{n}(x)\rangle_{f_{TT}}$ $\neq\frac{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\bm{k}_{\perp}|\bm{k}_{\perp}|^{n}f_{TT}(x,\bm{k}_{\perp}^{2})}{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\bm{k}_{\perp}f_{TT}(x,\bm{k}_{\perp}^{2})}\,,$ (36)
$\int_{0}^{1}\mathrm{d}x\langle k_{\perp}^{n}(x)\rangle_{h_{L}^{\perp}}$ $\neq\frac{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\bm{k}_{\perp}|\bm{k}_{\perp}|^{n}h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})}{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\bm{k}_{\perp}h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})}\,,$ (37)
$\langle k_{\perp}^{1,2}(x)\rangle_{g_{L}}$ is different, the minimum value is around $x=0.1$, the maximum values is around $x=1$, numerically,
$\langle k_{\perp}(x)\rangle_{g_{L}}$ $\in[0.28,0.42]\quad\text{GeV}\,,$ (38)
$\langle k_{\perp}^{2}(x)\rangle_{g_{L}}$ $\in[0.10,0.24]\quad\text{GeV},$ (39)
from Table 2, it is evident that although $\langle k_{\perp}^{1,2}(x)\rangle_{g_{L}}\neq\langle k_{\perp}^{1,2}(x)\rangle_{h_{L}^{\perp}}$, the equality holds for $\langle k_{\perp}^{n}\rangle_{g_{L}}$ and $\langle k_{\perp}^{n}\rangle_{h_{L}^{\perp}}$.
III.3 PDFs
Integrating the results of $\langle\Gamma\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp})$ over $\bm{k}_{\perp}$ one can obtain the four PDFs of a spin-1 target:
$\langle\gamma^{+}\rangle_{\bm{S}}^{(\lambda)}(x)$ $\equiv f(x)+S_{LL}f_{LL}(x)\,,$ (40a)
$\langle\gamma^{+}\gamma_{5}\rangle_{\bm{S}}^{(\lambda)}(x)$ $\equiv\lambda S_{L}g(x)\,,$ (40b)
$\langle\gamma^{+}\gamma^{i}\gamma_{5}\rangle_{\bm{S}}^{(\lambda)}(x)$ $\equiv\lambda S_{T}h(x).$ (40c)
Integrating Eq. (25), we can obtain
$f(x)$ $=\frac{N_{L}^{2}}{4\pi^{2}}\int\mathrm{d}\tau\frac{1}{\tau}(3x\bar{x}m_{\rho}^{2}+4\tau x^{2}\bar{x}^{2}m_{\rho}^{4})e^{-\tau B}$
$+\frac{N_{L}^{2}}{4\pi^{2}}\int\mathrm{d}\tau\frac{1}{\tau}(M^{2}+\frac{1}{\tau})e^{-\tau B}$
$+\frac{N_{T}^{2}}{4\pi^{2}}\int\mathrm{d}\tau(x^{2}+\bar{x}^{2})(\frac{1}{\tau}+\bar{x}xm_{\rho}^{2})e^{-\tau B}$
$+\frac{N_{T}^{2}}{4\pi^{2}}\int\mathrm{d}\tau(1+2x\bar{x}M^{2})e^{-\tau B}\,,$ (41a)
$g(x)$ $=\frac{3N_{T}^{2}}{8\pi^{2}}\int\mathrm{d}\tau e^{-\tau B}$
×$\times((2x-1)(\frac{1}{\tau}-(M^{2}-\bar{x}xm_{\rho}^{2}))+M^{2})\,,$ (41b)
$h(x)$ $=\frac{3N_{T}N_{L}M}{4\sqrt{2}\pi^{2}}\int\mathrm{d}\tau(\frac{1}{\tau}+2x\bar{x}m_{\rho}^{2})e^{-\tau B}\,,$ (41c)
$f_{LL}(x)$ $=\frac{3N_{L}^{2}}{4\pi^{2}}\int\mathrm{d}\tau\frac{1}{\tau}(3x\bar{x}m_{\rho}^{2}+4\tau x^{2}\bar{x}^{2}m_{\rho}^{4})e^{-\tau B}$
$=\frac{3N_{L}^{2}}{4\pi^{2}}\int\mathrm{d}\tau\frac{1}{\tau}(M^{2}+\frac{1}{\tau})e^{-\tau B}$
$-\frac{3N_{T}^{2}}{8\pi^{2}}\int\mathrm{d}\tau(x^{2}+\bar{x}^{2})(\frac{1}{\tau}+\bar{x}xm_{\rho}^{2})e^{-\tau B}$
$-\frac{3N_{T}^{2}}{8\pi^{2}}\int\mathrm{d}\tau(1+2x\bar{x}M^{2})e^{-\tau B}\,,$ (41d)
where $B=M^{2}+x(x-1)m_{\rho}^{2}$.
Figure 4: The upper panel: the unpolarized, tensor polarized, helicity and transversity PDFs of $\rho$ meson in this paper. The lower panel: the unpolarized, tensor polarized, helicity and transversity PDFs of $\rho$ meson in Ref. Ninomiya et al. (2017).
Note that the tensor polarized PDF $f_{LL}(x)$ is related to the single flavor distribution function $b_{1}(x)$,
$b_{1}(x)=\frac{1}{2}f_{LL}(x),$ (42)
$b_{1}(x)$ measures the difference in the spin projection of the $\rho$ meson, which depends solely on the quark-spin-averaged distribution
The PDFs of $\rho$ meson are depicted in Fig. 4. A comparison with Ref. Ninomiya et al. (2017) reveals that $f(x)$ and $f_{LL}(x)$ exhibit similarity, while the numerical value of our $h(x)$ is slightly smaller than their reported results. The primary disparity lies in the behavior of $g(x)$: in Ref. Ninomiya et al. (2017), $g(x)$ is negative for small values of $x$, whereas our findings indicate positivity across the entire range of $x\in[0,1]$.
The PDFs contain the information of the longitudinal momentum and the polarization carried by the partons. Here we define the $x$-moments of various PDFs as
$\langle x^{n}\rangle_{\alpha}=\int_{0}^{1}\mathrm{d}xx^{n}\alpha(x),$ (43)
where $\alpha(x)$ is the $\rho$ meson PDF.
The following sum rules should be satisfied,
$\int_{0}^{1}f(x)\mathrm{d}x=1,\quad\int_{0}^{1}f_{LL}(x)\mathrm{d}x=0,$ (44)
which means that in our model we have one valence $u$ quark and one valence $d$ quark in the $\rho$ meson. The sum rule of $f_{LL}$ corresponds to the $b_{1}$ sum rule in our model. The tensor polarized PDF $f_{LL}$ is symmetric around $x=1/2$, it is negative in the small and large $x$ region, and positive around the peak at $x=1/2$, $f_{LL}$ has two nodes at the model scale. $f(x)$ is also symmetric around $x=1/2$, we obtain the following results for the momentum sum rule of the valence quark:
$\int_{0}^{1}xf(x)\mathrm{d}x=\frac{1}{2},\quad\int_{0}^{1}xf_{LL}(x)\mathrm{d}x=0,$ (45)
this implies the valence quark and antiquark carry $100\%$ of the light-cone momentum, regardless of the spin quantization axis $S$ or the spin projection $\lambda$ of the hadron.
The spin sum rule of $g(x)$ is
$\Delta q\equiv\int_{0}^{1}g(x)\mathrm{d}x=0.505,$ (46)
which means the total contribution of the valence quark and antiquark to the spin of the $\rho$ meson is $50.5\%$, indicating that approximately $49.5\%$ of the contribution comes from the quark orbital angular momentum. In contrast, Ref. Ninomiya et al. (2017) reports this value as $56\%$, which is higher.
The sum rules of $h(x)$ is
$\int_{0}^{1}h(x)\mathrm{d}x=0.803,$ (47)
which indicates that $u$ quark of the $\rho$ meson’s tensor charge at the model scale is $0.803$, which is smaller than $0.94$ reported in Ref. Ninomiya et al. (2017). The values are considerably larger than the spin sum and approximate the naive quark model expectation of unity. In Table 3, we list the $\langle x^{0}\rangle$ and $\langle x^{1}\rangle$ moments of the various PDFs, and compared with the results in Refs. Ninomiya et al. (2017); Shi et al. (2022). Table  4 compares the $\langle x^{n}\rangle_{g}$ moments of the helicity PDF. It shows that for $\langle x^{0}\rangle_{g}$, the NJL result from the covariant method is close to the Lattice results. For $\langle x\rangle_{g}$, our result is a slightly larger than both the BSE and Lattice results. For $\langle x^{2}\rangle_{g}$, our result is larger than both the BSE and Lattice result, particularly the Lattice result.
Table 3: The $\langle x^{0}\rangle$ and $\langle x^{1}\rangle$ moments of $g(x)$ and $h(x)$ defined in Eq. (43). Here we compare with the results in the NJL model Ninomiya et al. (2017), Ref. Shi et al. (2022) employs the BSE-based LF-LFWFs and Ref. Kaur et al. (2024) utilizes LF quantization.
$\langle x^{0}\rangle_{g}$ $\langle x^{1}\rangle_{h}$
NJL  Ninomiya et al. (2017) 0.56 0.94
BSE  Shi et al. (2022) 0.67 0.79
LF quantization Kaur et al. (2024) 0.54 0.70
This work 0.505 0.803
Table 4: Comparison of $\langle x^{n}\rangle_{g}$ moments of the helicity PDF.
$\langle x^{0}\rangle_{g}$ $\langle x\rangle_{g}$ $\langle x^{2}\rangle_{g}$
NJL  Ninomiya et al. (2017) 0.558 0.369 0.251
BSE  Shi et al. (2022) 0.670 0.227 0.111
Lattice QCD Best et al. (1997) 0.570(32) 0.212(17) 0.077(34)
This work 0.505 0.279 0.183
III.4 Positivity constraints
From the probabilistic interpretation $\langle\Gamma\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})$, given by the quantities, one can obtain the inequalities:
$\langle\gamma^{+}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})\geqslant 0,$ (48)
$\langle\gamma^{+}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})\geqslant|\langle\gamma^{+}\gamma^{5}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})|,$ (49)
$\langle\gamma^{+}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})\geqslant|\langle\gamma^{+}\gamma^{i}\gamma^{5}\rangle_{\bm{S}}^{(\lambda)}(x,\bm{k}_{\perp}^{2})|.$ (50)
These relations must be valid for any spin quantization axis $\bm{S}$ and spin projection $\lambda$. Thus the spin-$1$ hadron, TMDs should satisfy the following relations  Ninomiya et al. (2017); Bacchetta and Mulders (2001)
$\text{I}.\quad f(x,\bm{k}_{\perp}^{2})\geqslant 0\,,$ (51a)
$\text{II}.\quad-\frac{3}{2}f(x,\bm{k}_{\perp}^{2})\leqslant f_{LL}(x,\bm{k}_{\perp}^{2})\leqslant 3f(x,\bm{k}_{\perp}^{2})\,,$ (51b)
$\text{III}.\quad|g_{L}(x,\bm{k}_{\perp}^{2})|\leqslant f(x,\bm{k}_{\perp}^{2})-\frac{1}{3}f_{LL}(x,\bm{k}_{\perp}^{2})\leqslant\frac{3}{2}f(x,\bm{k}_{\perp}^{2})\,,$ (51c)
$\text{IV}.\quad|h(x,\bm{k}_{\perp}^{2})|\leqslant f(x,\bm{k}_{\perp}^{2})+\frac{1}{6}f_{LL}(x,\bm{k}_{\perp}^{2})\leqslant\frac{3}{2}f(x,\bm{k}_{\perp}^{2}).$ (51d)
For the first constraint, the first diagram in Fig. 1 shows that our results satisfy the condition $f(x,\bm{k}_{\perp})\geqslant 0$. The remaining positivity constraints are illustrated in Fig. 5. These diagrams demonstrate that our $\rho$ meson TMDs fulfill the positivity constraints I–IV.
Figure 5: The positivity constraints for the quark TMDs of the $\rho$ meson expressed in Eq. (51).
Integrating TMDs equations above over $\bm{k}_{\perp}$ gives the following naive positivity conditions on the PDFs:
$\text{I}.\quad\quad f(x)\geqslant 0\,,$ (52a)
$\text{II}.\quad\quad-\frac{3}{2}f(x)\leqslant f_{LL}(x)\leqslant 3f(x)\,,$ (52b)
$\text{III}.\quad\quad|g_{L}(x)|\leqslant f(x)-\frac{1}{3}f_{LL}(x)\leqslant\frac{3}{2}f(x)\,,$ (52c)
$\text{IV}.\quad\quad|h(x)|\leqslant f(x)+\frac{1}{6}f_{LL}(x)\leqslant\frac{3}{2}f(x).$ (52d)
In Fig. 6, we show that our results satisfy the positivity constraints effectively. From Fig. 4, we observe that $f(x)>0$, indicating that the first condition is also met. Both Figs. 4 and 6 demonstrate that our $\rho$ meson PDFs satisfy constraints I–IV. In contrast to the findings in Ref. Ninomiya et al. (2017), our results reveal deviations from the relationships $f_{LL}=-\frac{3}{2}f$, $|g|=\frac{3}{2}f$ and $|h|=\frac{M}{m_{\rho}}\frac{3}{2}f$ at the endpoints $x=0$ and $x=1$, as shown in Fig. 6.
Figure 6: The positivity constraints for the quark PDFs of the $\rho$ meson expressed in Eq. (52).
III.5 Spin densities in the momentum space
The TMDs can be explained as the quark densities inside the hadron. The distribution of quark momentum inside the target can be defined by various polarization combinations represented through TMDs. Follow Ref. Kaur et al. (2021), the quark spin densities in the momentum space for the spin-1 target are defined as,
$\rho(x,k_{x},k_{y},(\lambda,\bm{\lambda}_{\perp}),(\Lambda,\bm{\Lambda}_{\perp}))=f+\lambda\Lambda g_{L}+\lambda\Lambda_{\perp}^{i}\frac{k_{\perp}^{i}}{m_{\rho}}g_{T}+\lambda_{\perp}^{i}\Lambda_{\perp}^{i}h+\lambda_{\perp}^{i}\Lambda\frac{k_{\perp}^{i}}{m_{\rho}}h_{L}^{\perp}$
$+(3\lambda^{2}-2)\left(\left(\frac{1}{6}-\frac{1}{2}\Lambda^{2}\right)f_{LL}+\Lambda\Lambda_{\perp}^{i}\frac{k_{\perp}^{i}}{m_{\rho}}f_{LT}+\left(\Lambda_{\perp}^{i}\Lambda_{\perp}^{j}-\frac{1}{2}\Lambda_{\perp}^{2}\delta_{ij}\right)\frac{k_{\perp}^{i}k_{\perp}^{j}}{m_{\rho}^{2}}f_{TT}\right),$ (53)
in the longitudinal direction, $\lambda$ and $\Lambda$ denote the spins of the quark and the target, respectively. These symbols represent different quantities from previous sections: $\lambda=\uparrow,\downarrow$ (or $+1$, $-1$) and $\Lambda=\uparrow,\downarrow$ (or $+1$, $-1$). The transverse spins of the quark and the target $\rho$ meson are denoted as $\bm{\lambda}_{\perp}=\Uparrow,\Downarrow$ (or $+1$, $-1$), and $\bm{\Lambda}_{\perp}=\Uparrow,\Downarrow$ (or $+1$, $-1$), respectively. In this paper, we consider transverse polarization along the $x$-direction.
To obtain all spin densities in the transverse momentum plane, we performed an integration of the longitudinal momentum fraction $x$. From Eq. (III.5) one can obtain $\rho_{\uparrow\uparrow}$ and $\rho_{\uparrow\downarrow}$
$\rho_{\uparrow\uparrow}(k_{x},k_{y})$ $=f+g_{L}-\frac{1}{3}f_{LL}\,,$ (54)
$\rho_{\uparrow\downarrow}(k_{x},k_{y})$ $=f-g_{L}-\frac{1}{3}f_{LL}\,,$ (55)
$\rho_{\downarrow\downarrow}=\rho_{\uparrow\uparrow}$ and $\rho_{\downarrow\uparrow}=\rho_{\uparrow\downarrow}$, therefore, our focus will be solely on studying $\rho_{\uparrow\uparrow}$ and $\rho_{\uparrow\downarrow}$. $\rho_{\uparrow\uparrow}$ represents the likelihood of locating the quark within the $\rho$ meson with its spin aligned to the spin of the composite system. $\rho_{\uparrow\uparrow}$ denotes the probability of both spins being in an anti-aligned state. The spin densities of the quarks in the momentum space for the $\rho$ meson are depicted in Fig. 7. In the first row, we have $\rho_{\uparrow\uparrow}$ and $\rho_{\uparrow\downarrow}$, both exhibiting an axially symmetric behavior in transverse momentum space. It is worth noting that $\rho_{\uparrow\uparrow}$ is larger than $\rho_{\uparrow\downarrow}$ due to the opposite signs of $f$ and $g_{L}$. Additionally, it is observed that while the covariant approach in Ref. Ninomiya et al. (2017) yields a negative value for $\rho_{\uparrow\downarrow}$, our calculation results in a positive value for this quantity.
Refer to captionRefer to caption
Figure 7: The $u$ quark spin densities of the $\rho$ meson in the momentum space: $\rho_{\uparrow\uparrow}(k_{x},k_{y})$ and $\rho_{\uparrow\downarrow}(k_{x},k_{y})$ in the first row; $\rho_{\Uparrow\Uparrow}(k_{x},k_{y})$ and $\rho_{\Uparrow\Downarrow}(k_{x},k_{y})$ in the second row; $\rho_{\uparrow\Uparrow}(k_{x},k_{y})$ and $\rho_{\uparrow\Downarrow}(k_{x},k_{y})$ in the third row; $\rho_{\Uparrow\uparrow}(k_{x},k_{y})$ and $\rho_{\Uparrow\downarrow}(k_{x},k_{y})$ in the fourth row;
$\rho_{\Uparrow\Uparrow}$ and $\rho_{\Uparrow,\Downarrow}$ are defined as
$\rho_{\Uparrow\Uparrow}(k_{x},k_{y})$ $=f+h+\frac{\bm{k}_{\perp}^{2}}{m_{\rho}^{2}}f_{TT}\,,$ (56)
$\rho_{\Uparrow\Downarrow}(k_{x},k_{y})$ $=f-h+\frac{\bm{k}_{\perp}^{2}}{m_{\rho}^{2}}f_{TT},$ (57)
and $\rho_{\Downarrow\Downarrow}=\rho_{\Uparrow\Uparrow}$ and $\rho_{\Downarrow\Uparrow}=\rho_{\Uparrow\Downarrow}$. $\rho_{\Uparrow\Uparrow}$ and $\rho_{\Uparrow\Downarrow}$ denote the combination of zero and two units of orbital angular momentum (OAM) transfer overlap terms. They are depicted in the second row of Fig. 7. It can be observed that they also exhibit axial symmetry in the transverse momentum space, with the numerical value of $\rho_{\Uparrow\Uparrow}$ being greater than that of $\rho_{\Uparrow\Downarrow}$.
$\rho_{\uparrow\Uparrow}$ and $\rho_{\uparrow\Downarrow}$ are defined as
$\rho_{\uparrow\Uparrow}(k_{x},k_{y})$ $=f+\frac{k_{x}}{m_{\rho}}g_{T}\,,$ (58)
$\rho_{\uparrow\Downarrow}(k_{x},k_{y})$ $=f-\frac{k_{x}}{m_{\rho}}g_{T}\,,$ (59)
$\rho_{\downarrow\Downarrow}=\rho_{\uparrow\Uparrow}$ and $\rho_{\downarrow\Uparrow}=\rho_{\uparrow\Downarrow}$. $\rho_{\uparrow\Uparrow}$ and $\rho_{\uparrow\Downarrow}$ represent the spin orientations of the quark and the $\rho$ meson, with longitudinal and transverse components, respectively. These are depicted in the third row of Fig. 7. In the momentum space, $\rho_{\uparrow\Uparrow}$ and $\rho_{\uparrow\Downarrow}$ exhibit distortion effects due to the term $\frac{k_{x}}{m_{\rho}}g_{T}$. Specifically, for $\rho_{\uparrow\Uparrow}$, the axis of symmetry shifts to $(k_{x}\simeq 0.09,k_{y}=0)$, while for $\rho_{\uparrow\Downarrow}$, it shifts to $(k_{x}\simeq-0.09,k_{y}=0)$.
$\rho_{\Uparrow\uparrow}$ and $\rho_{\Downarrow\downarrow}$ are defined as
$\rho_{\Uparrow\uparrow}(k_{x},k_{y})$ $=f+\frac{k_{x}}{m_{\rho}}h_{L}^{\perp}\,,$ (60)
$\rho_{\Uparrow\downarrow}(k_{x},k_{y})$ $=f-\frac{k_{x}}{m_{\rho}}h_{L}^{\perp}\,,$ (61)
$\rho_{\Downarrow\downarrow}=\rho_{\Uparrow\uparrow}$ and $\rho_{\Downarrow\uparrow}=\rho_{\Uparrow\downarrow}$. $\rho_{\Downarrow\downarrow}=\rho_{\Uparrow\uparrow}$ and $\rho_{\Downarrow\uparrow}=\rho_{\Uparrow\downarrow}$. The quantities $\rho_{\Uparrow\uparrow}$ and $\rho_{\Downarrow\downarrow}$ are depicted in the final row of Fig. 7, where they exhibit distortion effects due to the term $\frac{k_{x}}{m_{\rho}}h_{L}^{\perp}$. Specifically, for $\rho_{\Uparrow\uparrow}$, the axis of symmetry shifts to $(k_{x}\simeq-0.07,k_{y}=0)$, while for $\rho_{\Downarrow\downarrow}$, it moves to $(k_{x}\simeq 0.07,k_{y}=0)$. A comparison between these two rows reveals that the distortion of a longitudinally-polarized quark in a transversely-polarized target is opposite to that of a transversely-polarized quark in a longitudinally polarized target. This phenomenon arises from the fact that $g_{T}$ is positive and $h_{L}^{\perp}$ is negative.
IV Summary and conclusion
In this study, we assess the leading-twist transverse momentum dependent parton distributions (TMDs) of the $\rho$ meson using light-front wave functions (LFWFs) within the framework of the Nambu–Jona-Lasinio (NJL) model with proper time regularization.
We have observed that the TMDs $f(x,\bm{k}_{\perp}^{2})$, $g_{L}(x,\bm{k}_{\perp}^{2})$, $g_{T}(x,\bm{k}_{\perp}^{2})$, $h(x,\bm{k}_{\perp}^{2})$ and $f_{TT}(x,\bm{k}_{\perp}^{2})$ exhibit similar behavior, showing symmetry around $x=1/2$. As the transverse momentum $\bm{k}_{\perp}^{2}$ increases, these TMDs display a much weaker dependence on the light-cone momentum fraction $x$. This implies that a struck quark with large $\bm{k}_{\perp}^{2}$ has a much weaker preference for any particular value of $x$. For the longitudinal TMD $g_{L}(x,\bm{k}_{\perp}^{2})$, the maximum value shifts to occur at $x>1/2$. Conversely, for the transversely polarized TMD $h_{L}^{\perp}(x,\bm{k}_{\perp}^{2})$, the minimum value corresponds to a momentum fraction $x<1/2$. In terms of tensor polarized TMDs, $f_{LL}(x,\bm{k}_{\perp}^{2})$ and $f_{LT}(x,\bm{k}_{\perp}^{2})$, we observe that $f_{LL}(x,\bm{k}_{\perp}^{2})$ exhibits two peaks and is symmetric about $x=1/2$, while $f_{LT}(x,\bm{k}_{\perp}^{2})$ is antisymmetric and equals zero at $x=1/2$ for all $\bm{k}_{\perp}^{2}$.
We study the transverse momentum dependence of the $\rho$ meson TMDs, with a focus on their $\bm{k}_{\perp}$-weighted moments. In Ref. Ninomiya et al. (2017), the authors also calculate the $\rho$ meson TMDs in the NJL model using the covariant method. A comparison of $k_{\perp}$-weighted moments with their results shows that the numerical values are almost identical, except for $\langle k_{\perp}^{n}\rangle_{g_{L}}$, where our values are larger than theirs. Notably, our numerical values of $\langle k_{\perp}^{n}\rangle_{g_{L}}$ align with those of $\langle k_{\perp}^{n}\rangle_{h_{L}^{\perp}}$, consistent with findings in Refs. Kaur et al. (2021); Shi et al. (2022). Furthermore, we investigate the $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ of TMDs to determine typical transverse momenta of quark TMDs. Our results indicate that the typical transverse momenta of $\langle k_{\perp}(x)\rangle_{\alpha}$ fall within the range $[0.3,0.45]$ GeV and those of $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$ lie within $[0.15,0.30]$ GeV region.
The four $\rho$ meson PDFs can be obtained by integrating the TMDs over $\bm{k}_{\perp}$. A comparison with Ref. Ninomiya et al. (2017) reveals that $f(x)$, $f_{LL}(x)$, $h(x)$ and $h(x)$ are similar, but there is a difference in $g(x)$. In Ref. Ninomiya et al. (2017), $g(x)$ is negative when $x$ is small, whereas our results show that it is positive in the entire region of $x\in[0,1]$. We also conducted an investigation into the $x$-moments of various PDFs. Our spin sum rule indicates that $\Delta q=\langle x\rangle_{g}=0.505$, suggesting that the total valence quark and antiquark contribution to the spin of the $\rho$ meson is $50.5\%$, implying that about $49.5\%$ contribution comes from quark orbital angular momentum. The $x$-moments of $h(x)$ yield $\langle x\rangle_{h}=0.803$, indicating that the tensor charge at the model scale is 0.803.
The positivity constraints of $\rho$ meson TMDs and PDFs are investigated in this study. The relationships $f_{LL}=-\frac{3}{2}f$, $|g|=\frac{3}{2}f$, and $|h|=\frac{M}{m_{\rho}}\frac{3}{2}f$ at the endpoints $x=0$ and $x=1$, as reported in Ref. Ninomiya et al. (2017), do not hold in our results, which is evident from Fig. 6. The diagrams presented in Figs. 6 and 5 demonstrate that our findings align well with the positivity constraints.
Finally, we conducted a study on the spin densities within the $\rho$ meson in transverse momentum space. The spin densities are depicted in Fig. 7. From the diagram, it is evident that both the quark and the target exhibit axial symmetry when polarized in either longitudinal or transverse directions. However, when the quark is longitudinally polarized and the target is transversely polarized (or vice versa), a dipolar distortion is observed in the symmetric distribution. Furthermore, opposite distortions are observed in configurations where the quark and $\rho$ meson are longitudinally-transversely or transversely-longitudinally polarized. Another discovery pertains to the quark spin densities of the $\rho$ meson, specifically the $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)=f-\frac{1}{3}f_{LL}-g_{L}$. In the covariant approach, it is negative; however, $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)$ from LFWFs is positive.
For further investigation, our ongoing efforts to incorporate the effects of additional partons, such as sea quark contributions, may offer a more comprehensive understanding of the $\rho$ meson TMDs. Within the framework of the NJL model, these effects could be naturally represented by considering the influence of the virtual pion cloud surrounding the dressed quarks. Furthermore, we intend to account for higher Fock-state effects in future studies.
Acknowledgements.
Work supported by: the Scientific Research Foundation of Nanjing Institute of Technology (Grant No. YKJ202352), the Natural Science Foundation of Jiangsu Province (Grant No. BK20191472), and the China Postdoctoral Science Foundation (Grant No. 2022M721564).